Reports until 12:13, Tuesday 04 June 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 12:13, Tuesday 04 June 2024 (11626)
morning lab notes

-Achieve good laser lock with 775 light. Noticeable second beam on camera when locked. Not sure where its coming from but it should be diagnosed.

-Used plastic tool that comes in the tunable inductor kit to maximize the error signal amplitude by tuning the inductance. Error amp increased from 3mV pp to just over 5 mV pp. Note this still needs to be done on the 1550 path. Need alignement first though.

-Reprofiled the 775 light going to the caavity. MM is worse. May need to move some lenses.Note though that the power is very low going into the cavity which reduces the acuracy of using the beam profiler.

-Dramatically improved alignment for 775 using cavity mirrors. 0,0 is dominant enough to justify not moving lenses for now. Error amplitude in reased to ~9.5 mV. TRANS PD was misaligned as well.

-The 775 light can be seen on the SWIR camera. I used this for initial alignment by walking the 1550 beam to get close to the 775 beam that is on resonance. This was successful and we now have 1550 flashes in the cavity! The orientation to achieve this however is having the reflected beam go very high such that it is missing the mirrors that steer it to the REFL pd. This needs to be solved.